OpenStudy (anonymous):

Period? Asymptote equation? f(x)=cot4(x+pi/6)+3

OpenStudy (dumbcow):

standard period for tan/cot is pi period = pi/B = pi/4 asymptotes occur when dividing by zero cot = cos/sin find when sin[4(x+pi/6)] = 0 4(x+pi/6) = npi x + pi/6 = npi/4 x = npi/4 - pi/6